OpenStudy (anonymous):

find the product: (2x2-3x+5)(x-4) Sorry the exponents are big

OpenStudy (anonymous):

2x^3-11x^2+17x-20

OpenStudy (callisto):

(2x^2-3x+5)(x-4) = 2x^2 (x-4) -3x(x-4) +5(x-4) = 2x^3 - 2x^2 (4) - 3x(x) - 3x(-4) + 5x +5(-4) = 2x^3 - 8x^2 - 3x^2 + 12x + 5x -20 = Can you continue?

OpenStudy (anonymous):

is this accurate?

OpenStudy (anonymous):

Distributive property.\[(2x^2-3x+5)(x-4)=2x^2(x-4)-3x(x-4)+5(x-4)=2x^3-11x^2+17x-20\]

OpenStudy (anonymous):

thanks